Caffeine absorption.after caffeine is absorbed into the body 13% is eliminated from the body each hour.assume a person drinks an 8 Oz cup of brewed coffee containing 130mg of caffeine,and the caffeine is absorbed immediately into the body.Write a program to compute the following values.the number of hours required until 65mg(one-half the orginal amount remain in the body.the amount of caffeine in the body 24hours after the person drinks the coffee
looking for the code
1
Expert's answer
2012-04-24T11:05:56-0400
Just make the form in any VB version with 1 button, and add corresponding code.


The answer is 5 hours and 4mg.


Just in case, about 0,87:
(X*100%-X*13%)/100%=x*87%/100%=X*0,87


So to remove 13% we just multiply by 0,87





Public Class Form1

Dim caff_dose As Integer 'to count current dose
Dim hour As Integer 'to count hours
Dim caff_half_dose_time As Integer 'to hold found 1 question value
Dim hour24_left As Integer 'to hold found 2 question value


Private Sub Button1_Click(ByVal sender As System.Object, ByVal e As System.EventArgs) Handles Button1.Click

'initial values
caff_dose = 130
hour = 0


While caff_dose > 65 'the loop will end when the dose would fall below 65
hour = hour + 1 'increasing hours
caff_dose = caff_dose * 0.87 'decreasing dose
MsgBox("HALF DOSE TIME: on hour " & hour quot; we have " & caff_dose quot;mg") 'you can remove this line, just to show loop output
caff_half_dose_time = hour 'so when the loop stops we will have desired number
End While

'initial values again
caff_dose = 130
hour = 0


While hour < 24 'the loop will end on 24-th hour
hour = hour + 1 'same as above loop
caff_dose = caff_dose * 0.87 'same as above loop
MsgBox("24 HOURS: on hour " & hour quot; we have " & caff_dose quot;mg") 'you can remove this line, just to show loop output
hour24_left = caff_dose 'so we have the second value
End While


MsgBox("The number of hours required until 65mg amount remain in the body is " & caff_half_dose_time)
MsgBox("The amount of caffeine in the body 24hours after the person drinks the coffee is " & hour24_left)
End Sub
End Class